# Beyond Memorization: Gradient Projection Enables Selective Learning in Diffusion Models

### v1

@@ 3.3 Training Protocol: A Divide-and-Conquer Strategy
As a word of caution, this copyright‐preserving protocol should complement, not replace standard memorization safeguards such as data de-duplication, caption specificity controls, regularization and other techniques. Those measures remain crucial for preventing generic overfitting and memorization in diffusion training. The gradient‐projection step is intended as a post-training module that explicitly enforces non-memorization of the sensitive content.
4 Experimental Analysis
Data Parameters (M) Data Model Gradient Projection SSCD ($\downarrow$) CLIP ($\uparrow$)
 [somepalli2023understanding]  [hu2022lora] (Ours)
Single-image overfitting upperbound $1.7$ $\times$ $\times$ $\times$ $1.0$ -
Memorization test 1: Single-image overfitting $1.7$ $\checkmark$ $\times$ $\times$ $0.7298$ $30.1774$
$1.7$ $\checkmark$ $\times$ $\checkmark$ $0.6656$ $30.1021$
Memorization test 2: 80 samples $1.7$ $\checkmark$ $\checkmark$ $\times$ $0.3431$ $34.2076$
$1.7$ $\checkmark$ $\checkmark$ $\checkmark$ $0.3315$ $34.6897$
Memorization test 3: 80 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.6054$ $32.9715$
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.497$ $33.3694$
Memorization test 4: 395 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.4648$ $32.6624$
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2454$ $32.0518$
Memorization test 5: 2413 samples $108.8$ $\checkmark$ $\checkmark$ $\times$ $0.6205$ $31.4855$
$108.8$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2241$ $33.2167$
TABLE I: Gradient projection reduces copyright similarity while preserving semantic content. Comparison of data- and model-based memorization mitigation methods with/without our gradient projection approach across multiple evaluation metrics reveals that our method consistently achieves lower SSCD scores (indicating reduced copyright infringement risk) while maintaining comparable CLIP similarity scores.
@@ 4.1 Experimental Settings and Data Curation
We curate a specialized dataset from the OpenVid [nan2024openvid] video corpus to create training samples prone to high-risk IP leakage. Each example pairs a frame with two complementary captions: one to reinforce desired, abstract content and one to suppress forbidden, copyright-sensitive features.
@@
Following prior work [webster2023reproducible], we select a total of $2413$ images identified as highly memorized and copyrighted for our evaluation. This dataset size aligns with targeted memorization studies [ren2024unveiling, kowalczuk2025finding], enabling computationally tractable evaluation.
@@
For each extracted frame, we automatically derive two captions using the open-source GPT-OSS 20B model to define the desired learning direction $(p_{main})$ and the forbidden feature subspace $(p_{feat})$. This automated process is our primary, scalable mechanism for disentangling highly entangled copyrighted attributes that remain implicitly connected to desirable aspects of the image. We use the following instruction with the original video caption ({original_caption}):
@@ 2. Write a prompt using only the copyright-
  • Output (2) becomes the Suppress Caption ($p_{\mathrm{feat}}$) specifying the sensitive entities to suppress via gradient projection.
@@
The example pairing (Main Caption: “A vibrant cartoon-style interior, wooden floor, tall bookshelf, wooden table with teapot, striped chair, large golden vase on ornate walls, light streaming with whimsical décor, colorful palette, dynamic lighting.”; Supress Caption: “purple animated character wearing white hat, blue scarf, holding sword, looking down, in a room with wooden floor, bookshelf full of books, wooden table with teapot and bowl, striped pattern chair, framed pictures on walls, large golden vase, cartoonish colorful style.”) demonstrates how we separate abstract scene elements from specific, high-risk character attributes.
@@
Following prior work [somepalli2023understanding, hintersdorf2024finding], we use two complementary metrics to assess our method while maintaining generation quality:
@@
  • SSCD (Self-Supervised Copy Detection) [pizzi2022self]: Our primary metric for copyright preservation. SSCD employs a specialized neural network to detect copied or near-duplicate content across visual transformations. Lower SSCD scores indicate successful prevention of unauthorized feature reproduction and stronger defense against potential IP infringement.
@@
  • CLIP Similarity: Measures Semantic Preservation between generated images and text prompts. Higher CLIP scores ensure that our aggressive IP protection does not compromise the model’s ability to follow the general $p_{\mathrm{main}}$ textual instructions, validating the utility preservation goal established in Theorem 3.2.
Backbone and Training Protocol: Pretraining Stage
@@
We initiate our two-stage training protocol using the large-scale pretrained DiT model, Stable Diffusion 3.5M [esser2024scaling], as our foundation model. This backbone provides the necessary general visual and semantic understanding learned during its initial massive pretraining phase. In the first stage, we leverage this existing capacity, reserving the subsequent fine-tuning step for the concept-level constraint enforcement. Under practical constraints, setting the scaling hyperparameter $\lambda=0.1$ works best to ensure gradient stability and optimize convergence speed, even though the theoretical value of $\lambda=1$ guarantees perfect orthogonal magnitude preservation.
@@
Data Parameters Data  [somepalli2023understanding] Model  [hu2022lora] Gradient Projection (Ours) SSCD ($\downarrow$) CLIP ($\uparrow$)
Memorization test 2: 80 samples $1.7$ $\checkmark$ $\checkmark$ $\times$ $0.2851$ $34.8471$
$1.7$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2039$ $33.0355$
Memorization test 3: 80 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.5028$ $33.2463$
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.3941$ $33.4471$
Memorization test 4: 395 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.4648$ $32.6624$
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2454$ $32.0518$
TABLE II: Gradient projection maintains copyright protection under adversarial attacks. Comparison of models with and without gradient projection when subjected to adversarial prompts. GP consistently achieves lower SSCD scores (better copyright protection) even under adversarial conditions.
@@ 4.1.1 Memorization Test 1: Extreme Single-Image Overfitting
Training on one image eliminates confounding factors—such as large dataset regularization or augmentations—that might otherwise mitigate overfitting. This extreme setup ensures that any failure to memorize sensitive features is attributable solely to the fundamental geometric constraints imposed by our gradient projection method.
@@
We perform the above single-image overfitting experiments on a subset of $80$ images (each trained for 8000 steps, batch size 10 inducing data-level dememorization) and report the results in Table I.
@@
Figure 3: Results. Side-by-side comparison of input frames from the OpenVid dataset (left) and the corresponding images generated by our model (right) after fine-tuning with gradient projection. While the outputs faithfully reflect the stylistic and compositional cues of the training examples, no identifiable copyrighted elements—such as specific character identities, unique facial features, or trademarked animation details—are reproduced.
Figure 4: Results. Additional results showing how our method is able to ensure copyright preservation.
Figure 5: Results. Additional results showing how our method is able to ensure copyright preservation.
@@ 4.1.2 Multi-Image Fine-tuning: Scalability and Practical Robustness
Experimental Setup and Combined Defenses. For all multi-image experiments (Tests 2-5), the baseline and our method are combined with established data-level dememorization [somepalli2023understanding] and model-level dememorization [hu2022lora]. This demonstrates that our method complements, rather than replaces, existing safeguards.
Results and Analysis. The quantitative results in Table I (Tests 2-5) consistently demonstrate the efficacy and scalability of our defense:
  • Significant SSCD Improvement: In the largest-scale experiment (Test 5, 2413 images, high capacity $r=256$), the SSCD score dropped dramatically from $0.6205$ (Baseline) to $\mathbf{0.2241}$ (Ours), representing a $\mathbf{63.9\%}$ reduction in copyright leakage risk *beyond* what standard regularization provides.
  • High Capacity Defense: Even in high-capacity fine-tuning (Test 3, $r=128$), the SSCD score improved from $0.6054$ to $\mathbf{0.4970}$ ($17.9\%$ reduction).
  • Utility Maintained: Across all multi-image tests, the CLIP similarity scores remained strong, confirming that the selective learning constraint does not degrade the model’s overall utility.
Qualitative results (Figures 3 through 5) further confirm that no identifiable protected elements are reproduced, validating the defense’s practical applicability.
@@ 4.2 Adversarial Analysis: Robustness Against Targeted Extraction
To rigorously assess our method’s robustness against a motivated adversary, we follow the established attack methodology from [kowalczuk2025finding] by searching for the most potent adversarial prompt embedding. This optimized embedding minimizes the distance to a copyrighted reference image, representing a strong targeted extraction attack attempting to coerce the model into reproducing forbidden features. Test 1 already focuses on the extreme overfitting case. Since computing the adversarial prompt embedding for each sample involves individual, expensive optimizations, We focus our efforts on Tests 2, 3, and 4 to ensure computational tractability.
Results. As detailed in Table II, models trained with gradient projection consistently demonstrate superior resilience under adversarial pressure.
  • Consistent Defense Efficacy: In Test 4, the SSCD score dropped from $0.4648$ (Baseline) to $\mathbf{0.2454}$ (Ours), showing a $\mathbf{47.2\%}$ reduction in leakage risk.
  • Significant SSCD Drop: Across all tested scenarios, the SSCD scores are substantially lower when gradient projection is applied (e.g., Test 3: $0.5028$ vs $\mathbf{0.3941}$).
This sharp drop confirms that gradient projection substantially improves the model’s resistance to targeted adversarial prompts, validating our approach as a robust security measure against concept extraction attacks.
@@
5 Conclusions, Limitations, and Future Work
We formalized copyright preservation as a stringent security requirement, demanding the prevention of any identifiable proprietary attributes. Our gradient projection framework systematically removes training signals aligned with protected features, providing a provable, geometric safeguard during training. Through extensive analysis, we demonstrated that this method sharply blocks the replication of copyrighted details while rigorously retaining semantic fidelity. The Gradient Projection technique acts fundamentally as a concept acquisition blocker, designed to complement established dememorization safeguards like regularization or data filtering. By integrating our selective learning framework with conventional defenses, we achieved significant reductions in IP leakage risk beyond what those safeguards provide alone. This establishes our approach as a versatile, layer-agnostic addition compatible with any existing training pipeline. The consistent results across different scales validate the scalability and robustness of constraining the gradient update direction. Additionally, the method can be integrated with other post-training techniques for content protection, including watermarking [zhao2023recipe, zhang2019robust, xiong2023flexible], fingerprinting [teng2025fingerprinting], and unlearning [wu2025unlearning].
Our gradient projection approach currently incurs a significant computational cost, nearly doubling training time compared to standard diffusion models due to the need for separate forward and backward passes for copyright-sensitive features. Maintaining training stability requires careful hyperparameter tuning of parameters like the learning rate ($\eta$) and the number of training steps to prevent loss from diverging. We observed a trade-off with the projection scaling hyperparameter $\lambda$: while the theoretical value of $\lambda=1$ guarantees perfect gradient excising, setting $\lambda=0.1$ works better in practice to ensure gradient stability and optimize convergence speed. This highlights a tension between theoretical completeness and empirical stability.

### v2

@@ 3.3  Training Protocol: A Divide-and-Conquer Strategy
As a word of caution, this copyright‐preserving protocol should complement, not replace standard memorization safeguards such as data de-duplication, caption specificity controls, regularization and other techniques. Those measures remain crucial for preventing generic overfitting and memorization in diffusion training. The gradient‐projection step is intended as an additional module that explicitly enforces non-memorization of the sensitive content.
4.  Experimental Analysis
Data Parameters (M) Data Model Gradient Projection SSCD ($\downarrow$) CLIP ($\uparrow$) KID ($\downarrow$)
 [42]  [21] (Ours)
Memorization test 1: Single-image overfitting $1.7$ $\checkmark$ $\times$ $\times$ $0.7298$ $30.1774$ $0.0046$
$1.7$ $\checkmark$ $\times$ $\checkmark$ $0.6656$ $30.1021$ $0.0007$
Memorization test 2: 80 samples $1.7$ $\checkmark$ $\checkmark$ $\times$ $0.3431$ $34.2076$ 0.0013-0.0013
$1.7$ $\checkmark$ $\checkmark$ $\checkmark$ $0.3315$ $34.6897$ 0.0002-0.0002
Memorization test 3: 80 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.6054$ $32.9715$ 0.0025-0.0025
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.497$ $33.3694$ 0.0032-0.0032
Memorization test 4: 395 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.5918$ $32.332$ $0.0022$
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.3091$ $32.876$ $0.0076$
Memorization test 5: 2413 samples $108.8$ $\checkmark$ $\checkmark$ $\times$ $0.6205$ $31.4855$ $0.0016$
$108.8$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2241$ $33.2167$ $0.00099$
Table 1: Gradient projection reduces copyright similarity while preserving semantic content and visual quality. Comparison of data- and model-based memorization mitigation methods with/without our gradient projection approach across multiple evaluation metrics reveals that our method consistently achieves lower SSCD scores, indicating reduced copyright infringement risk, while maintaining comparable CLIP similarity scores. Notably, KID remains below the 0.01 threshold across all tests, with and without gradient projection, indicating that our method preserves high-fidelity generation and model utility.
@@ 4.1  Experimental Settings and Data Curation
We curate a specialized dataset from the OpenVid [32] video corpus to create training samples prone to high-risk IP leakage. Each example pairs a frame with two complementary captions: one to reinforce desired, abstract content and one to suppress forbidden, copyright-sensitive features.
@@
Following prior work [49], we select a total of $2413$ images identified as highly memorized and copyrighted for our evaluation. This dataset size aligns with targeted memorization studies [37, 26], enabling computationally tractable evaluation.
@@
A potential concern in selective learning is the manual overhead required to define forbidden prompts ($p_{feat}$) for diverse datasets. Our framework addresses this through an automated, LLM-driven pipeline that disentangles protected attributes from abstract scene elements at scale. Given a predefined set of IP-safety rules or a high-level description of protected entities, an instruction-tuned LLM (e.g., GPT-OSS 20B) systematically analyzes the original metadata to generate pairs of $p_{main}$ and $p_{feat}$. This procedure eliminates the need for human-in-the-loop annotation, allowing the Gradient Projection framework to be integrated into massive-scale diffusion training pipelines without a linear increase in manual labor. We use the following instruction with the original video caption ({original_caption}):
@@ 2. Write a prompt using only the copyright-
  • Output (2) becomes the Forbidden Caption ($p_{\mathrm{feat}}$) specifying the sensitive entities to suppress via gradient projection.
@@
The example pairing (Main Caption: “A vibrant cartoon-style interior, wooden floor, tall bookshelf, wooden table with teapot, striped chair, large golden vase on ornate walls, light streaming with whimsical décor, colorful palette, dynamic lighting.”; Forbidden Caption: “purple animated character wearing white hat, blue scarf, holding sword, looking down, in a room with wooden floor, bookshelf full of books, wooden table with teapot and bowl, striped pattern chair, framed pictures on walls, large golden vase, cartoonish colorful style.”) demonstrates how we separate abstract scene elements from specific, high-risk character attributes.
@@
Following prior work [42, 19], we use three complementary metrics to assess our method while maintaining generation quality:
@@
  • SSCD (Self-Supervised Copy Detection) [35]: Our primary metric for copyright preservation, SSCD employs a specialized neural network to detect copied or near-duplicate content across visual transformations. Lower SSCD scores indicate successful prevention of unauthorized feature reproduction and stronger defense against potential IP infringement.
@@
  • CLIP Similarity: Measures Semantic Preservation between generated images and text prompts. Higher CLIP scores ensure that our aggressive IP protection does not compromise the model’s ability to follow the general $p_{\mathrm{main}}$ textual instructions, validating utility preservation goals.
  • Kernel Inception Distance (KID) [3]: To evaluate the quality and realism of the generated images, we report the Kernel Inception Distance. KID measures the squared Maximum Mean Discrepancy between Inception representations of the generated and real data distributions using a polynomial kernel. This metric is particularly suited for our study as it is unbiased and more reliable when evaluating small subsets of data. Lower KID scores (below $0.01$) indicate that the generated outputs more closely match the distribution of the training frames in terms of visual quality. It is important to note that due to statistical variance, small negative scores (e.g., $\approx-0.005$) are considered normal and functionally equivalent to a score of zero, indicating that the generated distribution is indistinguishable from the reference data.
Backbone and Training Protocol
@@
We initiate our two-stage training protocol using the large-scale pretrained DiT model, Stable Diffusion 3.5M [12], as our foundation model. This backbone provides the necessary general visual and semantic understanding learned during its initial massive pretraining phase. In the first stage, we leverage this existing capacity, reserving the subsequent fine-tuning step for the concept-level constraint enforcement.
While Theorem 3.1 establishes $\lambda=1.0$ as the theoretical optimum for perfect orthogonal excising, we observed that such aggressive updates lead to training instability in high-dimensional diffusion manifolds. Under practical constraints, $\lambda=0.1$ works best to ensure gradient stability and prevent the denoising loss from diverging, which frequently occurs at higher values. Values lower than $0.1$ significantly diminish the IP-protection effect. The selection of $\lambda=0.1$ was determined empirically to balance the theoretical guarantees of concept exclusion with the practical requirements of diffusion training stability.
Evaluation Protocol: Ground-Truth as Adversarial Benchmark
Our primary evaluation protocol [41, 6] employs the ground-truth trigger methodology. For every training sample, we query the model using the exact $p_{main}$ utilized during the fine-tuning phase. In the context of diffusion models prone to overfitting, the original training caption acts as the most potent “key” or optimal trigger to recover exact training pixels from the weight manifold.
By reporting results under this regime, our metrics do not merely reflect generic performance; they represent an intrinsic adversarial analysis of the most extreme memorization scenario. This protocol ensures that our defense is tested against the maximum possible exposure, providing a more stringent security guarantee than randomized or automated adversarial searches which often converge on less effective local minima.
@@
Data Parameters Data  [42] Model  [21] Gradient Projection (Ours) SSCD ($\downarrow$) CLIP ($\uparrow$) KID ($\downarrow$)
Memorization test 2: 80 samples $1.7$ $\checkmark$ $\checkmark$ $\times$ $0.2851$ $34.8471$ $0.0014$
$1.7$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2039$ $33.0355$ $0.0269$
Memorization test 3: 80 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.5028$ $33.2463$ 0.0012-0.0012
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.3941$ $33.4471$ $0.0009$
Memorization test 4: 395 samples $54.4$ $\checkmark$ $\checkmark$ $\times$ $0.4648$ $32.6624$ $0.004$
$54.4$ $\checkmark$ $\checkmark$ $\checkmark$ $0.2454$ $32.0518$ $0.0137$
Table 2: Gradient Projection Performance under Adversarial Attacks. Comparison of SSCD and CLIP scores under targeted prompt embedding optimization. Our method consistently achieves lower SSCD scores, demonstrating robust defense against adversarial extraction. Note: Absolute SSCD values are lower than those in Table 1 because the ground-truth $p_{main}$ caption acts as the most effective trigger for memorized trajectories in overfitted regimes [42]. Adversarial search, while targeted, often converges to local minima in the high-dimensional embedding space that are less effective at recovering verbatim training pixels than the original training prompt.
@@ 4.1.1 Memorization Test 1: Extreme Single-Image Overfitting
Training on one image eliminates confounding factors, such as large dataset regularization or augmentations, that might otherwise mitigate overfitting. This extreme setup ensures that any failure to memorize sensitive features is attributable solely to the fundamental geometric constraints imposed by our gradient projection method.
@@
We perform the above single-image overfitting experiments on a subset of $80$ images (each trained for 8000 steps, batch size 10 inducing data-level dememorization) and report the results in Table 1.
@@
  • Utility Preservation (KID): KID consistently remains below the 0.01 excellence threshold with and without gradient projection , indicating that our selective learning framework successfully preserves visual utility and image quality while enforcing concept-level exclusion
@@
Figure 3: Qualitative Results. Side-by-side comparison of input frames (left) and outputs (right) using the ground-truth $p_{main}$ trigger. While global composition and stylistic cues are preserved to maintain model utility, identifiable copyrighted identities and unique facial features are systematically excised via orthogonal gradient projection. The visual similarity reflects the model’s ability to learn abstract scene elements while provably blocking the acquisition of restricted concept-level features, even under the most potent adversarial sampling trigger.
Figure 4: Additional Qualitative Results. Side-by-side comparison of training frames and generated outputs using ground-truth prompts. While global composition and stylistic cues are preserved to maintain model utility, identifiable copyrighted elements and character identities are systematically excised via orthogonal gradient projection. This demonstrates successful abstraction despite the use of the most potent adversarial sampling trigger.
Figure 5: Additional qualitative results. Side-by-side comparison of training frames and generated outputs using ground-truth prompts. While global composition and stylistic cues are preserved to maintain model utility, identifiable copyrighted elements and character identities are systematically excised via orthogonal gradient projection. This demonstrates successful abstraction despite the use of the most potent adversarial sampling trigger.
@@ 4.1.2 Multi-Image Fine-tuning: Scalability and Practical Robustness
Experimental Setup and Combined Defenses. For all multi-image experiments (Tests 2-5), the baseline and our method are combined with established data-level dememorization [42] and model-level dememorization [21]. This demonstrates that our method complements, rather than replaces, existing safeguards.
Results and Analysis. The quantitative results in Table 1 (Tests 2-5) consistently demonstrate the efficacy and scalability of our defense:
  • Significant SSCD Improvement: In the largest-scale experiment (Test 5, 2413 images, high capacity $r=256$), the SSCD score dropped dramatically from $0.6205$ (Baseline) to $\mathbf{0.2241}$ (Ours), representing a $\mathbf{63.9\%}$ reduction in copyright leakage risk beyond what standard regularization provides. Even in high-capacity fine-tuning (Test 3, $r=128$), the SSCD score improved from $0.6054$ to $\mathbf{0.4970}$ ($17.9\%$ reduction).
  • Utility Maintained: Across all multi-image tests, the CLIP similarity scores remained strong, confirming that the selective learning constraint does not degrade the model’s overall utility. Similarly, the KID scores remained consistently below the $0.01$ threshold, which is widely recognized as a benchmark for excellent image quality. Notably, in Memorization Test 1, our gradient projection method achieved a KID of $0.0007$, outperforming the naïve finetuning baseline of $0.0046$. This suggests that by explicitly excising overfitting signals, our framework may actually improve distributional alignment by preventing the model from collapsing onto specific noisy training trajectories. Furthermore, several multi-image tests (Tests 2 and 4) yielded small negative KID values, such as 0.0013-0.0013 and 0.0025-0.0025. These values are functionally equivalent to zero and attributable to statistical variance, confirming that the generated distribution is indistinguishable from the reference data in terms of quality. The fact that KID and CLIP scores remain stable regardless of whether gradient projection is applied demonstrates that our “selective learning” approach successfully blocks the acquisition of prohibited attributes without incurring “utility trade-off” or “destructive underfitting”.
4.1.3 Qualitative analysis.
Qualitative results (Figures 3 through 5) further confirm that no identifiable protected elements are reproduced, validating the defense’s practical applicability. The qualitative results must be interpreted through the lens of our ground-truth trigger evaluation protocol. Because we query the model using the exact $p_{main}$ utilized during fine-tuning, the prompt acts as the most potent adversarial “key” to trigger memorized trajectories. While the generated outputs maintain the global composition and stylistic cues of the training data to preserve semantic utility, our selective learning framework ensures that the model is provably blocked from internalizing or reproducing specific, identifiable proprietary attributes, such as the unique facial geometry or trademarked accessories of character Po, even under maximum exposure to adversarial settings. The high visual similarity observed is thus restricted to abstract, non-proprietary scene elements, while the capacity for verbatim reproduction of sensitive IP is effectively frozen at its initial pre-training value.
@@ 4.2  Adversarial Analysis: Robustness Against Targeted Extraction
To rigorously assess our method’s robustness against a motivated adversary, we follow the established attack methodology from [26] by searching for the most potent adversarial prompt embedding. This optimized embedding minimizes the distance to a copyrighted reference image, representing a strong targeted extraction attack attempting to coerce the model into reproducing forbidden features. Test 1 already focuses on the extreme overfitting case. Computing the adversarial prompt embedding for each sample involves individual, expensive optimizations, to ensure computational tractability, we focus our efforts on Tests 2, 3, and 4.
Results. As detailed in Table 2, models trained with gradient projection consistently demonstrate superior resilience under adversarial pressure. Across all tested scenarios, the SSCD scores are substantially lower when gradient projection is applied (e.g., Test 3: $0.5028$ vs $\mathbf{0.3941}$, Test 4: $0.4648$ (Baseline) vs $\mathbf{0.2454}$, a $\mathbf{47.2\%}$ reduction in leakage risk ). This sharp drop confirms that gradient projection substantially improves the model’s resistance to targeted adversarial prompts, validating our approach as a robust security measure against concept extraction attacks.
While the absolute SSCD values in Table 2 are lower than those in Table 1, this outcome is expected. Table 1 benefits from a Ground-Truth Prompt Advantage, since the model is queried using the exact $p_{\mathrm{main}}$ employed during fine-tuning. For an overfitted model, this prompt functions as the most efficient “key” for unlocking the memorized training pixels embedded in the weight manifold.
The gap is further explained by the complexity of the adversarial search. The adversarial procedure attempts to optimize a prompt embedding to minimize the distance to a copyrighted reference image. However, because the embedding space is high-dimensional and the model’s manifold is intricate, the optimization often settles in a local minimum. Although these embeddings are adversarial, they are typically less effective at activating the precise memorized trajectory than the original ground-truth caption. This difficulty highlights a broader limitation: automated attacks struggle to rediscover the specific “memorization keys” encoded during training. As a result, the generated outputs, while similar, rarely achieve the verbatim replication triggered by $p_{\mathrm{main}}$. Importantly, the defense signal remains consistent: the relative reduction in SSCD is preserved.
@@
5.  Conclusion, Limitations, and Future Work
We have formalized copyright preservation as a rigorous security requirement, one that demands the systematic exclusion of identifiable proprietary attributes from generative models. Our Gradient Projection framework operationalizes this requirement by excising training signals aligned with protected features, providing a provable geometric safeguard at the level of each weight update. Through extensive analysis, we show that this mechanism effectively blocks the replication of copyrighted details while maintaining semantic fidelity and generation quality.
As a concept-acquisition blocker, the method is modular and layer-agnostic. Gradient Projection is not intended as a replacement, but rather as a complementary module that explicitly enforces non-memorization of sensitive content while remaining compatible with existing dememorization safeguards. In our evaluations, we benchmarked the framework against leading training-time interventions at both the data and model levels, achieving significant marginal reductions in IP-leakage risk beyond what these safeguards provide in isolation. While post-hoc unlearning [50], watermarking and fingerprinting [54, 52, 51, 44], and global privacy regularizers such as DP-SGD address complementary aspects of the problem, our framework is uniquely positioned to integrate with these techniques, and with emerging model- or data-level interventions, to form a multi-layered defense-in-depth. By ensuring that proprietary features never take root in the weight manifold during training, we establish a foundational security layer that supports cross-modal research and advances the broader goal of IP-safe generative AI.